The MPC in your economy is 0.70 with an autonomous consumption of Tk 80 billion. Your economy is a open economy with I, T, R, G and XN with an exogenous amount of 20, 50, 20, 60 and 5 billion Taka, respectively. Find out your equilibrium GDP with two, three and four sector economy. Do you think your GDP will increase for:

TK 10 billion extra transfer payment;
Tk 10 billion new government purchases ; and
Tk 5 billion increase in the net export
If all of the above are considered together ?
1
Expert's answer
2012-02-28T07:37:45-0500
In open economy GDP: Y = C+I+G+Xn.C - consumption: C = m(DI) + b,
where: m = marginal propensity to consume, DI = disposable income, b = autonomous consumption (expenditure).
DI = Y - T, where Y is income (or GDP) and T represents total tax revenues.
Y = 0.7*(Y - 50) + 80 + 20 + 60 + 5.
Y = 53.8 billion.

New government purchases and increase in the net export will cause the increase in GDP according to expenditure approach of its calculation.
The above used method is expenditure approach.
For income approach, we must exclude transfer payments. Their effect is inversely proportional. Business transfer payments can be thought of as an extra cost tacked onto the price of goods, over and above factor payments, that compensates for bad debts incurred by the household sector.
Income approach:
GDP = Compensation of employees + Rent + Interest + Proprietor’s Income + Corporate Profits + Indirect business taxes + Depreciation + Net foreign factor income